OpenStudy (anonymous):

Any one with any complicated Indefinite integrals/3-D questions?

OpenStudy (nincompoop):

\[\int\limits \sqrt{\tan x \cos^2x} dx\]

terenzreignz (terenzreignz):

$$\large \int x^x dx$$

OpenStudy (anonymous):

thanks peeps

OpenStudy (anonymous):

but the second one, go kill yourself :|